### v1

@@ 6.2 Empirical Behavior of the Number of Pieces in $C_{t}$
This observation can be attributed to two key design mechanisms in our algorithm:
@@
    Range-based pruning: The function $C_{t}(I_{i}^{t})$ models the cost of ending day $t$ with inventory level $I_{i}^{t}$. As established in Section 4.4, the feasible domain for $I_{i}^{t}$ is $[0,U_{i}-d_{i}^{t}]$, where $U_{i}$ is the maximum inventory and $d_{i}^{t}$ is the demand on day $t$. This domain applies to both transitions (17) and (18) in the DP recursion. Before performing convolutions or minimizations, we discard any segments of the piecewise linear functions that fall entirely outside this interval. Additionally, segments that do not cover any integer-valued inventory levels are also removed, as they contain no feasible solutions. These two steps significantly reduce the number of pieces carried forward at each stage.
@@ 6.3 Performance of HGS-IRP on IRP with penalized stock-outs
Figure 7: Trade-off between stock-out cost and routing/inventory cost for a specific instance (35 customers and 5 vehicles) under varying stock-out penalties $\rho$, which varies from 0 to 300 in integer increments and decreases from left to right.
Figure 7 illustrates the trade-off observed on a representative instance when varying the penalty parameter $\rho$. The total demand in this case is 2478. The $x$-axis shows the stock-out quantity, and the total quantity delivered to retailers is displayed on top of each column. The $y$-axis represents the sum of the routing and inventory costs, with routing costs shown in black and inventory costs shown as shaded segments. The plot shows that lower values of $\rho$ result in higher stock-out quantities and reduced total cost. In particular, routing costs decrease consistently as stock-outs increase, as the algorithm avoids serving some retailers when the penalty is low. This leads to a shorter or less frequent delivery schedule. Conversely, as $\rho$ increases, the algorithm favors full delivery, reducing stock-outs but increasing overall cost due to more extensive routing and higher inventory levels. Note that multiple values of $\rho$ can yield the same best delivery decision. As a result, several distinct $\rho$ values may correspond to a single bar in the figure, with identical stock-out quantities and cost decompositions. It is observed that once $\rho$ exceeds 185, no stock-out events occur. Consequently, for $\rho\geq 186$, all corresponding bars in the figure coincide with the first bar, as increasing the penalty further only reinforces stock-out avoidance but does not change the outcome.
@@
Appendix A (Table 3) reports results for all tested instances under the four values of $\rho$. Instances are grouped by number of retailers, inventory cost type (HC and LC), and time horizon. For each configuration, we present the best solution value, the average solution value over 10 runs, and the corresponding computational time. These results confirm a general trend: higher $\rho$ values typically result in increased total cost and runtime. HGS-IRP maintains stable performance across the range of penalty settings, suggesting robustness in handling various stock-out scenarios.
@@ 7 Conclusions
The operator is implemented through a dynamic programming (DP) procedure, using piecewise-linear state functions and supported by preprocessing and acceleration techniques to ensure practical efficiency. As demonstrated in our experiments, this design enables the calculation of reinsertion schedules within microseconds, making it useful for systematic solution refinement within metaheuristics. We integrated the operator into the Hybrid Genetic Search (HGS) framework, where it serves as the sole component responsible for modifying delivery schedules and quantities. The resulting HGS-IRP algorithm achieves state-of-the-art performance: it discovers new best-known solutions for 365 out of 1,040 benchmark instances and improves the best feasible solutions for 191 out of 270 challenging larger-scale IRP instances introduced in 2023.
The research perspectives stemming from this work are numerous. While the integration of the DS operator into HGS already yields good performance —leveraging HGS’s effectiveness across a wide range of VRP variants— the operator itself is generic and can be embedded into other metaheuristics more specifically tailored to the IRP. Moreover, this search strategy could be adapted to other deterministic IRP variants, as well as to more complex settings such as two-stage stochastic IRPs and dynamic IRPs, which are increasingly relevant in real-world applications. In these contexts, the operator could be adapted to solve sample average approximation-based deterministic approximations of stochastic problems, either as a standalone approach within a rolling horizon framework, or as a generator of solutions for training imitation learning models, such as the one developed in Greif et al. (2024).
@@ Acknowledgments
The authors thank Dr. Khalid Laaziri for his assistance in IT-related matters, including but not limited to resolving server issues and problems with Python and Matplotlib, and Dr. Brett Kraabel for help with writing.

### v2

@@ 6.2 Empirical Behavior of the Number of Pieces in $C_{t}$
Two key design mechanisms in our algorithm explain this observation:
@@
    Range-based pruning: The function $C_{t}(I_{i}^{t})$ models the cost of ending day $t$ with inventory level $I_{i}^{t}$. As established in Section 4.4, the feasible domain for $I_{i}^{t}$ is $0,U_{i}-d_{i}^{t}]$, where $U_{i}$ is the maximum inventory and $d_{i}^{t}$ is the demand on day $t$. This domain applies to both transitions (17) and (18) in the DP recursion. Before performing convolutions or minimizations, we discard any segments of the piecewise linear functions that fall entirely outside this interval. Additionally, segments that do not cover any integer-valued inventory levels are also removed, as they contain no feasible solutions. These two steps significantly reduce the number of pieces carried forward at each stage.
@@ 6.3 Performance of HGS-IRP on IRP with penalized stock-outs
Figure 7 illustrates the trade-off observed on a representative instance when varying the penalty parameter $\rho$. The $x$-axis shows the stock-out quantity, and the total quantity delivered to retailers is displayed on top of each column. The $y$-axis represents the sum of the routing and inventory costs, with routing costs shown in black and inventory costs shown as shaded segments. Each bar also reports the total quantity delivered to all customers, which decreases as the stock-out amount increases. The plot shows that lower values of $\rho$ result in higher stock-out quantities and reduced total cost. In particular, routing costs decrease consistently as stock-outs increase, as the algorithm avoids serving some retailers when the penalty is low. This leads to a shorter or less frequent delivery schedule. Conversely, as $\rho$ increases, the algorithm favors full delivery, reducing stock-outs but increasing overall cost due to more extensive routing and higher inventory levels. Note that multiple values of $\rho$ can yield the same best delivery decision. As a result, several distinct $\rho$ values may correspond to a single bar in the figure, with identical stock-out quantities and cost decompositions. It is observed that once $\rho$ exceeds 91, no stock-out events occur. Consequently, for $\rho\geq 92$, all corresponding bars in the figure coincide with the first bar, as increasing the penalty further only reinforces stock-out avoidance but does not change the outcome.
@@
Figure 7: Trade-off between stock-out amount and routing/inventory cost for a specific instance (15 customers and 5 vehicles) under varying stock-out penalties $\rho$, which varies from 1 to 300 in integer increments.
@@
Table 3 also reports detailed results for all tested instances under the four values of $\rho$. Instances are grouped by number of retailers, inventory cost type (HC and LC), and time horizon. For each configuration, we present the best solution value, the average solution value over 10 runs, and the corresponding computational time. These results confirm a general trend: higher $\rho$ values typically result in increased total cost and runtime. We additionally report the gap (Gap) in average cost across $\rho$ values, using the $\rho=1,000,000$ case as a benchmark which approximates the no-stock-out case. This highlights the impact of stock-out penalties on the overall cost structure. As expected, higher $\rho$ values yield higher total costs, as the routing plans adjust to avoid larger stock-out penalties. Under the HC (high holding cost) setting, cost differences across $\rho$ are smaller, since high holding and stock-out costs already discourage stock-outs. By contrast, in the LC setting, small $\rho$ values produce relatively low routing-cost solutions, as stock-outs remain cheaper than replenishment. Overall, HGS-IRP maintains stable performance across the range of penalty settings, suggesting robustness in handling various stock-out scenarios.
Table 3: Best results (Best), average results (Avg), average runtime in seconds (T(s)) on instances of varying sizes for different stock-out penalty settings.
Data
Set
$n$ $\rho=50$ $\rho=100$ $\rho=300$ $\rho=1,000,000$
Best Avg Best Avg Best Avg Best Avg
Best T(s) Avg Gap(%) Best T(s) Avg Gap(%) Best T(s) Avg Gap(%) Best T(s) Avg
LC
 HH=3
5 587.08 16.9 587.08 -71.78 948.97 24.8 948.97 -54.38 1674.99 48.9 1674.99 -19.48 2080.31 87.1 2080.31
10 1497.86 44.5 1497.86 -50.94 2264.81 59.7 2264.81 -25.82 2955.57 108.0 2955.57 -3.20 3053.13 139.0 3053.13
15 1921.61 68.7 1921.61 -38.37 2540.49 87.7 2540.49 -18.52 3031.66 130.9 3031.97 -2.75 3117.85 155.4 3117.85
20 2553.68 88.6 2553.68 -30.75 3212.50 121.1 3212.50 -12.88 3611.01 162.5 3611.01 -2.07 3687.48 184.5 3687.49
25 3124.38 115.2 3126.72 -20.64 3643.10 159.4 3643.22 -7.53 3909.73 211.2 3914.53 -0.64 3936.38 215.9 3939.80
30 3353.37 149.2 3353.42 -15.05 3777.17 198.3 3777.38 -4.32 3920.38 253.3 3921.61 -0.66 3947.63 274.8 3947.74
35 3664.29 163.8 3668.62 -12.60 4088.36 239.6 4089.93 -2.57 4187.51 286.1 4187.64 -0.24 4197.61 297.4 4197.70
40 3782.98 205.9 3795.75 -12.20 4102.82 262.3 4107.49 -4.99 4271.98 339.1 4272.86 -1.17 4315.37 387.7 4323.25
45 4022.20 227.4 4028.15 -9.97 4293.90 293.0 4299.02 -3.92 4458.13 361.2 4463.54 -0.24 4471.83 386.8 4474.28
50 4676.77 262.4 4684.09 -10.59 5020.13 381.7 5033.01 -3.93 5181.11 485.6 5191.22 -0.91 5232.76 539.9 5239.03
Avg 2918.42 134.2 2921.70 -23.24 3389.23 182.8 3391.68 -10.89 3720.21 238.7 3722.49 -2.20 3804.04 266.8 3806.06
LC
HH=6
5 2114.42 84.3 2114.42 -64.15 3318.14 174.0 3318.15 -43.75 5192.22 372.3 5192.30 -11.98 5898.73 621.4 5898.77
10 3944.96 169.9 3944.96 -47.59 5732.51 324.8 5732.56 -23.84 7302.65 797.4 7304.50 -2.95 7524.83 1016.5 7526.86
15 5469.63 310.0 5469.87 -30.48 6937.18 618.4 6938.69 -11.82 7728.22 1028.9 7734.11 -1.71 7858.97 1189.9 7868.57
20 7038.84 470.3 7040.12 -25.78 8654.39 973.2 8660.29 -8.71 9370.41 1563.2 9390.04 -1.01 9475.69 1680.2 9486.06
25 8044.29 733.5 8048.87 -20.23 9382.32 1414.2 9402.07 -6.82 9983.37 1770.7 10006.16 -0.84 10063.19 1865.5 10090.52
30 8635.57 1187.4 8644.47 -13.02 9628.74 1707.7 9646.79 -2.94 9902.33 2046.3 9932.44 -0.06 9910.42 2072.6 9938.70
35 9594.55 1470.9 9617.88 -11.37 10512.11 1937.7 10541.70 -2.86 10813.82 2142.5 10841.46 -0.09 10819.22 2182.5 10851.70
40 9695.89 1640.3 9750.94 -11.38 10597.52 2053.2 10634.63 -3.35 10940.88 2202.3 10988.62 -0.13 10953.99 2225.9 11003.23
45 10441.67 1951.6 10466.43 -9.36 11176.42 2138.1 11209.45 -2.92 11500.24 2248.4 11542.45 -0.04 11509.90 2262.3 11547.06
50 12418.36 2141.8 12474.57 -7.78 13171.02 2279.9 13224.56 -2.23 13420.02 2366.0 13484.94 -0.31 13460.64 2368.9 13526.78
Avg 7739.82 1016.0 7757.25 -20.63 8911.04 1362.1 8930.89 -8.62 9615.42 1653.8 9641.70 -1.35 9747.56 1748.6 9773.82
HC
HH=3
5 2635.88 63.7 2635.88 -4.55 2733.03 81.2 2733.48 -1.01 2754.16 83.1 2754.16 -0.26 2761.47 83.7 2761.47
10 4829.84 125.3 4830.24 -0.54 4846.23 139.8 4846.28 -0.21 4856.50 145.6 4856.56 0.00 4856.50 143.1 4856.58
15 5491.10 144.4 5491.72 -0.71 5509.68 159.7 5509.68 -0.38 5524.66 165.1 5524.72 -0.11 5530.87 170.2 5530.88
20 7066.76 170.8 7067.11 -0.36 7082.51 180.1 7082.51 -0.14 7090.65 183.2 7090.65 -0.03 7092.57 181.9 7092.57
25 8421.57 204.3 8421.96 -0.10 8428.26 215.7 8428.45 -0.02 8429.46 228.0 8429.98 0.00 8429.50 221.9 8429.97
30 9656.95 270.2 9657.25 -0.04 9658.73 293.2 9658.84 -0.02 9660.57 301.5 9660.59 0.00 9660.57 295.7 9660.69
35 10166.57 348.7 10167.74 -0.03 10167.88 378.4 10168.28 -0.03 10170.59 369.8 10171.19 0.00 10170.63 360.3 10171.22
40 10919.59 408.2 10933.05 -0.17 10932.19 438.6 10952.13 0.01 10934.26 463.1 10953.13 0.02 10935.95 462.8 10951.37
45 11958.89 391.3 11964.93 -0.04 11961.90 419.3 11965.66 -0.04 11966.25 430.9 11977.74 0.06 11966.17 420.6 11970.11
50 13519.58 573.5 13529.78 -0.03 13522.07 635.8 13531.65 -0.02 13523.54 639.6 13530.39 -0.03 13523.45 609.7 13534.32
Avg 8466.67 270.0 8469.97 -0.31 8484.25 294.2 8487.70 -0.10 8491.06 301.0 8494.91 -0.01 8492.77 295.0 8495.92
HC
HH=6
5 7313.49 487.1 7313.82 -0.79 7352.60 570.9 7352.92 -0.26 7367.91 643.0 7368.72 -0.05 7371.85 601.3 7372.14
10 10849.74 911.9 10854.04 -0.80 10905.93 1041.9 10910.54 -0.28 10931.97 1069.5 10937.92 -0.03 10935.99 1060.2 10941.57
15 13066.26 1256.9 13074.88 -0.32 13104.46 1297.8 13116.45 -0.01 13106.68 1396.8 13117.23 0.00 13106.99 1323.3 13117.17
20 16307.98 1634.5 16321.56 -0.09 16319.36 1692.4 16335.14 -0.01 16321.38 1736.8 16336.75 0.00 16322.96 1688.1 16336.76
25 18320.33 1905.2 18350.60 -0.18 18340.80 1842.6 18370.67 -0.07 18358.87 1884.7 18382.77 0.00 18358.73 1865.3 18383.15
30 20810.58 2016.5 20835.51 0.01 20810.16 2040.6 20837.67 0.02 20808.01 2088.0 20836.47 0.02 20805.62 2028.9 20833.26
35 22607.50 2103.9 22643.27 0.02 22607.88 2161.8 22644.03 0.02 22614.14 2149.1 22644.51 0.02 22606.38 2072.9 22639.84
40 24146.43 2029.4 24203.13 0.07 24144.45 2127.1 24227.38 0.17 24143.91 2111.3 24224.99 0.16 24147.37 2179.5 24186.13
45 27129.80 617.0 27129.80 0.22 27129.80 617.0 27129.80 0.22 27070.40 559.0 27165.30 0.35 27036.50 2400.0 27071.16
50 30397.51 2282.3 30491.50 0.10 30413.11 2277.1 30518.17 0.19 30425.21 2275.0 30497.33 0.12 30405.67 2361.1 30460.95
Avg 19094.96 1524.5 19121.81 -0.06 19112.85 1566.9 19144.28 0.05 19114.85 1591.3 19151.20 0.09 19109.81 1758.0 19134.21
HC 50 30615.71 7200.0 30697.45 -0.05 30619.20 7200.0 30696.18 -0.06 30618.82 7200.0 30699.39 -0.05 30629.73 7200.0 30713.76
100 51806.43 7200.2 51970.67 -0.01 51806.30 7200.2 51959.48 -0.03 51803.57 7200.2 51953.48 -0.05 51803.26 7200.2 51976.93
200 95497.49 7200.9 95754.60 0.00 95523.94 7201.0 95763.75 0.01 95534.84 7201.0 95753.63 0.00 95526.64 7201.0 95757.36
Avg 59306.54 7200.4 59474.24 -0.01 59316.48 7200.4 59473.14 -0.02 59319.08 7200.4 59468.83 -0.02 59319.88 7200.4 59482.68
LC 50 12551.19 7200.0 12615.56 -8.06 13336.96 7200.0 13405.90 -2.30 13623.71 7200.0 13689.28 -0.23 13660.14 7200.0 13721.51
100 16341.00 7200.1 16444.51 -2.51 16656.30 7200.2 16754.31 -0.68 16783.32 7200.2 16867.82 0.00 16787.39 7200.2 16868.29
200 23910.12 7200.9 24110.79 -1.11 24135.77 7200.9 24347.91 -0.14 24183.88 7201.0 24389.36 0.03 24176.11 7200.9 24382.03
Avg 17600.77 7200.4 17723.62 -3.28 18043.01 7200.4 18169.37 -0.84 18196.97 7200.4 18315.49 -0.05 18207.88 7200.4 18323.94
@@ 7 Conclusions
The operator is implemented through a dynamic programming (DP) procedure, using piecewise-linear functions and supported by preprocessing and acceleration techniques to ensure practical efficiency. It is designed to be metaheuristic-independent: it only requires an incumbent solution and can therefore be embedded as a large-scale neighborhood within various frameworks, such as Iterated Local Search (ILS), adaptive Large Neighborhood Search (LNS), or other metaheuristics. As demonstrated in our experiments, this design enables the calculation of reinsertion schedules within microseconds, making it useful for systematic solution refinement within metaheuristics. We integrated the operator into the Hybrid Genetic Search (HGS) framework, where it serves as the sole component responsible for modifying delivery schedules and quantities. The resulting HGS-IRP algorithm achieves state-of-the-art performance: it discovers new best-known solutions for 365 out of 1,040 benchmark instances and improves the best feasible solutions for 191 out of 270 challenging larger-scale IRP instances introduced in 2023.
The research perspectives stemming from this work are numerous. While the integration of the DS operator into HGS already yields good performance, the operator itself is generic and can be embedded into other methods more specifically tailored to the IRP. Moreover, it could be extended to richer IRP variants, such as those with time windows, heterogeneous fleets, multiple depots, or to the production routing problem, by adapting the evaluation function to incorporate the corresponding constraints and cost components. Moreover, this search strategy could be adapted to other deterministic IRP variants, as well as to more complex settings such as two-stage stochastic IRPs and dynamic IRPs, which are increasingly relevant in real-world applications. In these contexts, the operator could be adapted to solve sample average approximation-based deterministic approximations of stochastic problems, either as a standalone approach within a rolling horizon framework or as a generator of solutions for training imitation learning models, such as the one developed in greif2024combinatorial.
@@ Acknowledgments
The authors thank the three anonymous reviewers whose comments helped to improve and clarify this manuscript. They also thank Dr. Khalid Laaziri for his assistance in IT-related matters, including but not limited to resolving server issues and problems with Python and Matplotlib, and Dr. Brett Kraabel for help with writing. This work was supported by SCALE-AI through its Research Chairs program, the National Natural Science Foundation of China (72571234), and the SRIBD Grant (J00220250002).
{APPENDICES}{APPENDICES}
8 Considering Possible Backorders
The IRP formulation considered in this paper assumes lost sales: that is, unmet demand in a given period is not carried over and instead incurs an immediate penalty cost. This reflects many real-world applications, such as retail, vending machines, or seasonal products, where backlogging is either impractical or undesirable. Nevertheless, from an algorithmic perspective, supporting backorders requires only minimal modifications to the DS operator and no modification of HGS. In particular, this requires enlarging the DP state space and adapting the cost-update function.
State Space. In the current model, the inventory level at the start of each period $I_{i}^{t}$ is restricted to the range $0,U_{i}]$, where $U_{i}$ is the retailer’s capacity. To allow backorders, this domain should be expanded to $B_{\max},U_{i}]$, where $B_{\max}$ is the maximum backlog allowed for retailer $i$ (e.g., the sum of all demands). This change increases the number of DP states and thus raises the computational complexity of the operator from $O(U_{i}KH)$ to $O((U_{i}+B_{\max})KH)$ per retailer.
Cost Function Modification. The rest of the DP structure, including the transition functions $f_{1}$ and $f_{2}$, remains unchanged. The only modification occurs in the final cost function update, where the stockout penalty is replaced by a per-period backlog cost. Specifically, we define:
$C_{t}(I_{i}^{t})=\hat{C}_{t}(I_{i}^{t})+p_{i}\cdot\max(0,-I_{i}^{t}),$ (22)
where $p_{i}$ is the backlog penalty per unit per day for retailer $i$, and $\hat{C}_{t}$ is the continuation cost computed recursively. The term $\max(0,-I_{i}^{t})$ captures the amount of backlog at the start of period $t$.
Backtracking. In the backtracking phase, the delivery quantity $q_{i}^{t}$ must satisfy both the current demand and any carried-over backlog. If $I_{i}^{t-1}<0$, the algorithm will naturally allocate enough delivery in period $t$ to reduce or clear the backlog. This behavior is seamlessly supported by the modified state space and cost function, requiring no further structural change. This simple extension allows the DS operator to balance routing, inventory holding, and backlog costs, making it applicable to contexts where backorders are acceptable or even necessary.